suppose that two objects attract each other with a gravitational force of 16units. if the distance between the two objects is doubled, what is the new force of attraction between the two objects?


1
Expert's answer
2021-11-01T12:53:42-0400
"F_{new}=\dfrac{Gm_1m_2}{(2d)^2}=\dfrac{Gm_1m_2}{4d^2}=\dfrac{1}{4}\dfrac{Gm_1m_2}{d^2}=\dfrac{F}{4},""F_{new}=\dfrac{16\ units}{4}=4\ units."